OpenStudy (__meggaan):

For his chemistry experiment, Martin needs 3 liters of a 40% alcohol solution. The lab has two containers, one with 20% alcohol solution and the other with 50% alcohol solution. And I need to find a set of linear equations. please help me

OpenStudy (__meggaan):

@Frozen15

OpenStudy (__meggaan):

@AloneS

OpenStudy (frozen15):

what do you think? Any ideas, thoughts?

OpenStudy (__meggaan):

Well, I started to figure out the problem and when I plugged my answers back into the equation, it was wrong

OpenStudy (frozen15):

Hmm can you tell/show me your equation you did?

OpenStudy (__meggaan):

so, I had to identify the variables which were x=20% alcohol solution and y=50%solution. then x+y=3 but I don't think thats right. I feel like that's probably where I screwed up. I was just following a problem I had that was similar

OpenStudy (frozen15):

Hmm i feel you did do right here but of course we have to MAKE SURE. But you got 3 for your answer? im a little confsued

OpenStudy (__meggaan):

x+y=3 .20+.50=1.2 (because I multiplied 3*.40) .20x+.50y=1.2 .20(3-y) + .50y = 1.2 .6-.20y+.50y=1.2 .6-.7y=1.2 -.6 -.6 -.7y=.6 /-.7 y=-.857142857 x=3-y x=3-(-.857142857) x=3.857142857

OpenStudy (__meggaan):

@Frozen15

OpenStudy (__meggaan):

@AloneS

OpenStudy (__meggaan):

I just need help :.(..